find the regression analysis, what is the value of a,b, and what is expected y if x=18

X: 17 12 10 20

Y: 10 10 8 12


1
Expert's answer
2021-10-11T05:10:14-0400
"\begin{matrix}
 & X & Y & XY & X^2 & Y^2 \\
 & 17 & 10 & 170 & 289 & 100 \\
 & 12 & 10 & 120 & 144 & 100 \\
 & 10 & 8 & 80 & 100 & 64 \\
 & 20 & 12 & 240 & 400 & 144 \\
 Sum =& 59 & 40 & 610 & 933 & 408 \\
\end{matrix}"

"\bar{X}=\dfrac{1}{n}\displaystyle\sum_{i=1}^nX_i=\dfrac{59}{4}=14.75"

"\bar{Y}=\dfrac{1}{n}\displaystyle\sum_{i=1}^nY_i=\dfrac{40}{4}=10"

"SS_{XX}=\displaystyle\sum_{i=1}^nX_i^2-\dfrac{1}{n}(\displaystyle\sum_{i=1}^nX_i)^2"

"=933-\dfrac{(59)^2}{4}=62.75"

"SS_{YY}=\displaystyle\sum_{i=1}^nY_i^2-\dfrac{1}{n}(\displaystyle\sum_{i=1}^nY_i)^2"

"=408-\dfrac{(40)^2}{4}=8"

"SS_{XY}=\displaystyle\sum_{i=1}^nX_iY_i-\dfrac{1}{n}(\displaystyle\sum_{i=1}^nX_i)(\displaystyle\sum_{i=1}^nY_i)"

"=610-\dfrac{59(40)}{4}=20"

The regression coefficients (the slope "m," and the y-intercept  "n") are obtained as follows:


"m=\dfrac{SS_{XY}}{SS_{XX}}=\dfrac{20}{62.75}=0.318725"

"n=\bar{Y}-m\bar{X}=10-\dfrac{20}{62.75}(14.75)=5.298805"

We find that the regression equation is:


"Y=5.298805+0.318725X"



Correlation coefficient


"r=\dfrac{SS_{XY}}{\sqrt{SS_{XX}}\sqrt{SS_{YY}}}=\dfrac{20}{\sqrt{62.75}\sqrt{8}}=0.892644"

Strong positive correlation.


"r^2=\dfrac{(20)^2}{62.75(8)}=0.796813"

"Y(18)=5.298805+0.318725(18)=11.035855"